OpenStudy (anonymous):

Is the following always, sometimes, or never true?m + 0 = 0

OpenStudy (lgbasallote):

anything added to 0 is the number itself i.e 0 + 0 = 0 1 + 0 = 1 2 + 0 = 2 3 + 0 = 3 and so on

OpenStudy (anonymous):

so if m is 0, what would m+0 be? and if m is something else, say 4, what would m+0 be? Answer those questions and you should be able to figure out the correct response.